OpenStudy (anonymous):

Rewrite with positive exponents. Assume that even roots are of nonnegative quantities and that all denominators are nonzero. (6xy/y)^-3

OpenStudy (anonymous):

heres a rule: x^-1=1/x

OpenStudy (anonymous):

im lost on this one

OpenStudy (anonymous):

basically on this problem well just flip the fraction to make the exponent positive

OpenStudy (anonymous):

try to simplify 6xy/y=?

hartnn (hartnn):

is the question \(\huge (\frac{6xy}{y})^{-3}\) then u can cancel out the y in numerator and denominator and write that as \(\huge ({6x}{})^{-3}\)

hartnn (hartnn):

that has negative exponent, to make it positive , we would use \(\huge a^{-n}=\frac{1}{a^n}\) to get \(\huge (6x)^{-3}=\frac{1}{(6x)^3}\) did u get this @RenaeMonroe

OpenStudy (anonymous):

yes

hartnn (hartnn):

good, then that would be your final answer.

OpenStudy (anonymous):

6x^3

hartnn (hartnn):

noes 1/(6x)^3 \(\huge \frac{1}{(6x)^3}\)

OpenStudy (anonymous):

thats the final answer 1/(6x)^3

hartnn (hartnn):

yup.

OpenStudy (anonymous):

how simple thank you so much

hartnn (hartnn):

welcome :)